If the two acute angles of a right angeled triangle are in the ratio7:8,find these angles
Answer by fcabanski(1391)

The sum of the two acute angles of a right triangle is 90.


7x + 8x = 90


15x = 90


x = 6


The angles are 7*6 = 42 degrees and 8*6 = 48 degrees